A 3000-lb. car has an engine which can deliver 80 hp to the rear wheels. What is the maximum velocity at which the car can climb a 15 degree hill?

1
Expert's answer
2022-06-07T17:37:05-0400

Given:

"3000\:\rm lb=13344\:N"

"80\:\rm hp=58840\: W"


"P=Fv\cos\theta"

"v=\frac{P}{F\cos\theta}=\frac{58840}{13344*\cos 15^\circ}=4.6\:\rm m/s"